OpenStudy (anonymous):

write the parametric representation of \(\large \rm z=4x^2+y^2\)

OpenStudy (anonymous):

is it z=4t x=2sin t y=4cos t ?

OpenStudy (tkhunny):

Well, let's see... Take t = 0 z=4t = 4(0) = 0 x=2sin t = 2(0) = 0 y=4cos t = 4(1) = 4 0 = 4(0^2) + 4^2 = 0+16 = 16 <== Well, THAT doesn't seem right!

OpenStudy (anonymous):

oh God :( so whats the right presentation ?

OpenStudy (anonymous):

@rational

OpenStudy (rational):

x = 1/2cost y = sint wont this work ?

OpenStudy (anonymous):

and z= ?

OpenStudy (rational):

x=2sin t y=4cos t also is fine right ? im not sure what you are parameterizing it for... z = 1 i guess

OpenStudy (anonymous):

ermm 

OpenStudy (rational):

since its a surface we need two variables may be... idk im at temporary memory loss :s

OpenStudy (rational):

Alright, it looks bit complicated http://tutorial.math.lamar.edu/Classes/CalcIII/ParametricSurfaces.aspx

OpenStudy (anonymous):